25 September UT

Afternoon

  • Andrea changed the software to fix the problem of single-chip observations leaving the detector in a bad bias state. Now integrations which follow single-chip images do not crash.
  • Fernando and Vincenzo are swapping back in the original red controller (ID #2). Last night we had the temperature problem with ID #1. In fact they found that the red controller fan is running slow and Fernando said that the actual temperature was increasing steadily. Original controller was tested all day in the lab.
  • 18:00 Original controller back in place (ID#2). Not working - cryostat needs to be dismounted and brought to lab. Planned for morning.
  • 19:45 start observing with Blue.
  • 21:00 stop for wind.

-- OlgaKuhn - 25 Sep 2008
Topic revision: r3 - 25 Sep 2008, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback